OpenStudy (anonymous):

Which ratio forms a proportion with 16 over 28? A. 4 OVER 7 B.1 OVER 2 C.32 OVER 48 D.9 OVER 12

OpenStudy (anonymous):

16/28=4*4/7*4=4/7